# PRIMT: Preference-based Reinforcement Learning with Multimodal Feedback and Trajectory Synthesis from Foundation Models

### v1

@@ 1 Introduction
Reinforcement learning (RL) has shown great success in various robotics domains [1, 2, 3, 4], yet it remains reliant on carefully designed reward functions. In many practical scenarios, designing an informative reward function is highly challenging, as task objectives are often implicit and multi-faceted [5]. Preference-based RL (PbRL)  [6, 7, 8] has emerged as a promising alternative to address this challenge by learning reward models from human comparative feedback over robot trajectories, providing a more intuitive means of aligning robotic systems with human intent [9, 10, 11, 12]. Nevertheless, the extensive human input required for preference labeling restricts the scalability of PbRL [13].
@@
To mitigate this bottleneck, recent work has explored leveraging foundation models (FMs), e.g., large language models (LLMs) and vision-language models (VLMs), as synthetic feedback sources, drawing on their broad world knowledge [14, 15, 16, 17]. Compared to using FMs to design dense reward functions [18, 19, 20] or provide auxiliary contrastive signals [21, 22, 23, 24], incorporating them as evaluators within PbRL offers a potentially more efficient and robust paradigm.
@@
However, obtaining reliable and high-quality FM feedback remains challenging, primarily due to the dominant reliance on single-modality evaluation. LLM-based approaches [15, 16] interpret structured textual projections of trajectories, such as sequential arrays of state-action pairs, enabling sophisticated temporal reasoning over procedural logic and motion progression [25, 26]. However, these textual descriptions can be abstract or incomplete, making LLMs prone to hallucinations of key events, especially when inferring fine-grained spatial interactions [27, 28]. On the other hand, VLM-based methods [14] analyze spatial cues from visual renderings of robot trajectories, such as final-state images or intermediate frames, effectively capturing spatial goal completion [29]. Yet these methods often overlook subtle temporal dynamics within the trajectory [30, 31]. Consequently, relying on either modality alone risks incomplete or unreliable feedback (see Appendix A for more analysis), highlighting the need for a more comprehensive multimodal evaluation framework.
@@
Furthermore, even if feedback from FMs reaches human-expert-level quality, PbRL still faces two intrinsic challenges: i) Query ambiguity: trajectory pairs often exhibit uniformly low quality in early training stages. This happens when they are generated from random or weakly optimized policies and lack task-relevant variations, making it hard to elicit meaningful preferences [7, 32]; and ii) Credit assignment: even when reliable trajectory-level preferences are available, it often remains difficult to accurately attribute the observed preference differences to specific states or actions [33, 34, 35]. Without effective state-action-level credit assignment, the learned reward model may result in misaligned behaviors in subsequent RL training [33]. Meanwhile, FMs have shown strong abilities in planning [36, 37], control [38, 39], and causal reasoning [40, 41]. These advances lead us to the following question: Can FMs move beyond serving as passive preference providers to be actively leveraged to mitigate query ambiguity and improve credit assignment in PbRL?
@@
i) Multimodal feedback fusion, which enhances synthetic feedback quality by combining the complementary advantages of LLM- and VLM-based evaluations. Rather than directly feeding multimodal trajectory representations into generic multimodal FMs, PRIMT adopts a hierarchical neuro-symbolic preference fusion strategy. It first performs intra-modal fusion to produce modality-specific labels and confidence estimations. Then, inter-modal fusion is conducted using probabilistic soft logic (PSL) [42], which infers the final preference label via structured and interpretable reasoning over multimodal evaluation outputs and trajectory context, enabling robust aggregation of preference beliefs from heterogeneous sources.
ii) Bidirectional trajectory synthesis, which leverages FMs to actively enhance reward learning in PbRL. In the foresight phase, LLMs generate diverse task-aligned trajectories to initialize the trajectory buffer. Unlike prior work [16, 36] that assumes FM-generated trajectories are optimal, we treat them as semantically meaningful anchors for informative comparisons, reducing early-stage query ambiguity. In the hindsight phase, LLMs are prompted to generate counterfactual trajectories via causal reasoning based on the structural causal model (SCM) [43]. When a clear preference is detected, the model identifies causal steps and applies minimal edits to reverse the preference, producing counterfactuals that highlight critical distinctions. To better exploit these samples for credit assignment, we introduce a causal-aware auxiliary loss that enforces reward separation at edited steps while ensuring consistency elsewhere. This enables more precise preference attribution, thereby improving the efficiency and generalization of the learned reward model in downstream RL training.
@@
  • We conduct extensive experiments across 2 locomotion and 6 manipulation tasks from the DMC [44], MetaWorld [45], and ManiSkill [46] benchmarks, demonstrating that PRIMT consistently outperforms state-of-the-art baselines. Ablation studies provide insight into component effectiveness, and we further validate PRIMT’s real-world applicability on a Kinova Jaco robot.
@@ 2 Related Works and Preliminaries
Foundation models refer to large-scale pre-trained models with strong generalization and reasoning capabilities across tasks [47]. Recent work has explored leveraging FMs to address the reward engineering challenge in RL. One line of research uses coding LLMs to directly generate executable reward functions [18, 19, 20]. Another employs VLMs as contrastive reward signals [21, 22, 23, 24]; for example, RoboCLIP [24] rewards agents by aligning trajectory images with task descriptions or demonstrations. However, such explicit FM-based reward signals are often noisy and high-variance [14]. Recent approaches have adopted the PbRL paradigm, using FMs as synthetic evaluators to generate trajectory-level preference labels and train reward models: PrefCLM [15] and RL-SaLLM-F [16] leverage LLMs to analyze numerical state-action sequences, while RL-VLM-F [14] uses VLMs to assess final-state images of robot trajectories. These approaches have shown improved performance over FM-generated scalar rewards. Our work builds on this PbRL-with-FM direction but introduces two key innovations. First, PRIMT adopts a multimodal evaluation scheme that combines VLM and LLM perspectives via hierarchical neuro-symbolic fusion, improving the robustness and quality of synthetic supervision. Second, rather than using FMs solely for passive evaluation, we actively incorporate them to facilitate reward learning via trajectory synthesis.
@@
PbRL aims to learn a reward model $r_{\psi}$ from human comparative feedback over pairs of robot trajectories [7]. A trajectory $\sigma$ is defined as a sequence of states and actions $\{(s_{1},a_{1})\ldots,(s_{T},a_{T})\}$ with a length of $T$. The annotator provides a preference label $\Upsilon\in\{-1,0,1\}$ for each pair $(\sigma^{A},\sigma^{B})$, where $\Upsilon=1$ indicates that $\sigma^{A}$ is preferred, 0 means $\sigma^{B}$ is preferred, and 1-1 denotes indecision. A preference predictor is constructed using the Bradley-Terry model [48] to estimate the preference probabilities. The likelihood that $\sigma^{A}$ is preferred over $\sigma^{B}$ is computed as:
@@
PbRL relies on informative preference queries to train effective reward models. However, trajectory pairs often exhibit low task-relevant diversity, leading to query ambiguity [8]. This ambiguity is especially pronounced in early training stages, when trajectories are uniformly low-quality and incoherent due to randomly initialized policies [49]. Prior works address this by selecting maximally distinguishable or uncertain samples [32, 50, 8] or initializing the reward model with expert demonstrations [51, 52]. More recent works employ LLMs to revise ambiguous trajectories during training, assuming that the edited versions are task-complete and preferred [16]. In contrast, we proactively initialize the replay buffer with LLM-generated trajectories that are diverse and task-aligned. Unlike prior work, we do not assume these are optimal, but use them as preference anchors to support more informative and efficient early-stage evaluation.
@@
Another core challenge in PbRL is the granularity mismatch between trajectory-level preference supervision and the desired state-action-level reward signal [33]. This mismatch introduces uncertainty in attributing credit to specific decisions, impairing both the alignment and generalization of the learned reward model [53]. Prior work mitigates this issue by training transformer-based world models that estimate state importance [33, 54, 55], or by collecting additional human annotations to highlight key moments [34]. In contrast, our approach requires neither extra supervision nor architectural changes. Inspired by causal counterfactual reasoning [56, 43, 57, 58], we prompt LLMs to generate hindsight-based counterfactual trajectories by minimally editing key decision points in the preferred trajectory to reverse the preference. By asking, “What minimal change would make this trajectory less preferred?”, we obtain contrastive examples that expose the underlying reasons for preference. To effectively leverage these counterfactuals in reward learning, we introduce a causal-aware auxiliary loss. It enforces reward separation at edited points while maintaining consistency in the unedited parts, leading to more precise credit assignment.
@@ 3.1 Multimodal Feedback Generation and Fusion
Given a trajectory pair $(\sigma^{A},\sigma^{B})$ sampled from the trajectory buffer, we first obtain their textual projections $text(\sigma^{A})$ and $text(\sigma^{B})$ for LLM-based evaluation, following [16]. These projections organize each trajectory into dimension-specific sequences, capturing structured temporal patterns across state and action components in a format that enhances semantic interpretability. For VLM-based evaluation, instead of using all frames or final-state images as in prior work [14], we propose a hybrid keyframe extraction method to capture both low-level motion cues and high-level behavior transitions while avoiding visual overload: i) near-zero velocity detection identifies frames where the robot motion is minimal, typically marking subgoal completions or transitional pauses [59]; ii) smoothing residual peaks detect high-curvature or abrupt motion transitions by comparing the raw trajectory to its smoothed version, capturing key motion shifts [60]; and iii) change point detection segments the trajectory into semantically coherent phases to identify structural high-level task changes [61]. We take the union of the selected frames from all three methods, together with the first and last steps of the trajectory, to form the final keyframe sets $kvis(\sigma^{A})$ and $kvis(\sigma^{B})$. Full details are provided in Appendix B.
@@
To efficiently model these latent dependency structures among inputs, heuristics, and decisions, we employ Probabilistic Soft Logic (PSL) [42], a probabilistic framework representing entities of interest as logical atoms interconnected by weighted first-order logic rules. Specifically, we define four rules to guide inter-modal preference fusion:
@@ 3.2 Bidirectional Trajectory Synthesis
Prior to PbRL training, we employ LLMs to generate bootstrapped trajectories that exhibit diverse, semantically meaningful, and task-aligned behaviors, providing a warm-start initialization for the trajectory buffer. Inspired by structured code-generation paradigms [38, 36], we adopt a three-step CoT strategy: i) generate a high-level, multi-step action plan from the task specification; ii) translate each step into executable code snippets that implement concrete motion primitives; and iii) execute these programs under varied initial conditions (e.g., robot start positions) and strategy parameters (e.g., height to approach the target) to synthesize a diverse set of plausible trajectories. Compared to directly prompting LLMs to generate low-level trajectory arrays [16], our method improves physical feasibility and semantic coherence by grounding trajectory synthesis in program logic. The generated trajectories are considered as bootstrapped demonstrations rather than optimal ones, subsequently evaluated by our multimodal feedback module. Combined with strategic sampling schemes, such as uncertainty-based sampling [8], these trajectories serve as informative preference anchors when paired with exploration trajectories, reducing ambiguity in early-stage preference queries and accelerating reward learning.
@@
During PbRL training, whenever a clear preference is identified by the multimodal feedback module, we prompt LLMs to perform hindsight reasoning to generate counterfactual variants of the preferred trajectory. This process follows a three-step reasoning pattern based on the structural causal model [43]:
@@
ii) Action: Select a key step $t^{*}\in T^{*}$ for minimal intervention, generating a counterfactual trajectory $\sigma^{*}_{cf}$ that reverses the original preference. This involves modifying some critical state-action features at the selected step, such as introducing a small gripper delay or adding a local end-effector position perturbation. The rest of the trajectory remains identical to the original while we allow the LLMs to apply light smoothing to the immediate neighbors (e.g., 2-3 steps before and after the intervention) to ensure physical continuity and avoid abrupt state transitions. Multiple counterfactual variants can be generated through repeated LLM sampling, providing a diverse set of sub-preferred alternatives. Following the minimal edit principle [62, 63], we filter the generated counterfactuals based on the L1 distance between the edited state-action pairs and the original, ensuring a small deviation threshold.

### v2

@@ 1 Introduction
Reinforcement learning (RL) has shown great success in various robotics domains [wang2022deep, wang2024initial, kober2013reinforcement, jo2024cognitive], yet it remains reliant on carefully designed reward functions. In many practical scenarios, designing an informative reward function is highly challenging, as task objectives are often implicit and multi-faceted [gupta2022unpacking]. Preference-based RL (PbRL)  [christiano2017deep, lee2021pebble, lee2021b] has emerged as a promising alternative to address this challenge by learning reward models from human comparative feedback over robot trajectories, providing a more intuitive means of aligning robotic systems with human intent [bai2022training, wang2024personalization, wang2022skill, wang2022feedback]. Nevertheless, the extensive human input required for preference labeling restricts the scalability of PbRL [park2022surf].
@@
To mitigate this bottleneck, recent work has explored leveraging foundation models (FMs), e.g., large language models (LLMs) and vision-language models (VLMs), as synthetic feedback sources, drawing on their broad world knowledge [wang2024rl, wang2025prefclm, tu2024online, lee2023rlaif]. Compared to using FMs to design dense reward functions [maeureka, venutocode, xietext2reward] or provide auxiliary contrastive signals [ma2023liv, mahmoudieh2022zero, rocamonde2023vision, sontakke2023roboclip], incorporating them as evaluators within PbRL offers a potentially more efficient and robust paradigm.
@@
However, obtaining reliable and high-quality FM feedback remains challenging, primarily due to the dominant reliance on single-modality evaluation. LLM-based approaches [wang2025prefclm, tu2024online] interpret structured textual projections of trajectories, such as sequential arrays of state-action pairs, enabling sophisticated temporal reasoning over procedural logic and motion progression [liu2024st, xiong2024large]. However, these textual descriptions can be abstract or incomplete, making LLMs prone to hallucinations of key events, especially when inferring fine-grained spatial interactions [li2024advancing, zhang2025mitigating]. On the other hand, VLM-based methods [wang2024rl] analyze spatial cues from visual renderings of robot trajectories, such as final-state images or intermediate frames, effectively capturing spatial goal completion [luo2024vision]. Yet these methods often overlook subtle temporal dynamics within the trajectory [ding2024language, bossen2025can]. Consequently, relying on either modality alone risks incomplete or unreliable feedback (see Appendix A for more analysis), highlighting the need for a more comprehensive multimodal evaluation framework.
@@
Furthermore, even if feedback from FMs reaches human-expert-level quality, PbRL still faces two intrinsic challenges: i) Query ambiguity: trajectory pairs often exhibit uniformly low quality in early training stages. This happens when they are generated from random or weakly optimized policies and lack task-relevant variations, making it hard to elicit meaningful preferences [lee2021pebble, feng2024comparing]; and ii) Credit assignment: even when reliable trajectory-level preferences are available, it often remains difficult to accurately attribute the observed preference differences to specific states or actions [vermahindsight, holk2024polite, liang2021reward]. Without effective state-action-level credit assignment, the learned reward model may result in misaligned behaviors in subsequent RL training [vermahindsight]. Meanwhile, FMs have shown strong abilities in planning [kwon2024language, ahn2022can], control [liang2023code, huang2023voxposer], and causal reasoning [kiciman2023causal, di2023towards]. These advances lead us to the following question: Can FMs move beyond serving as passive preference providers to be actively leveraged to mitigate query ambiguity and improve credit assignment in PbRL?
@@
i) Multimodal feedback fusion, which enhances synthetic feedback quality by combining the complementary advantages of LLM- and VLM-based evaluations. Rather than directly feeding multimodal trajectory representations into generic multimodal FMs, PRIMT adopts a hierarchical neuro-symbolic preference fusion strategy. It first performs intra-modal fusion to produce modality-specific labels and confidence estimations. Then, inter-modal fusion is conducted using probabilistic soft logic (PSL) [bach2017hinge], which infers the final preference label via structured and interpretable reasoning over multimodal evaluation outputs and trajectory context, enabling robust aggregation of preference beliefs from heterogeneous sources.
ii) Bidirectional trajectory synthesis, which leverages FMs to actively enhance reward learning in PbRL. In the foresight phase, LLMs generate diverse task-aligned trajectories to initialize the trajectory buffer. Unlike prior work [tu2024online, kwon2024language] that assumes FM-generated trajectories are optimal, we treat them as semantically meaningful anchors for informative comparisons, reducing early-stage query ambiguity. In the hindsight phase, LLMs are prompted to generate counterfactual trajectories via causal reasoning based on the structural causal model (SCM) [pearl2010causal]. When a clear preference is detected, the model identifies causal steps and applies minimal edits to reverse the preference, producing counterfactuals that highlight critical distinctions. To better exploit these samples for credit assignment, we introduce a causal-aware auxiliary loss that enforces reward separation at edited steps while ensuring consistency elsewhere. This enables more precise preference attribution, thereby improving the efficiency and generalization of the learned reward model in downstream RL training.
@@
  • We conduct extensive experiments across 2 locomotion and 6 manipulation tasks from the DMC [tassa2018deepmind], MetaWorld [yu2020meta], and ManiSkill [gu2023maniskill2] benchmarks, demonstrating that PRIMT consistently outperforms state-of-the-art baselines. Ablation studies provide insight into component effectiveness, and we further validate PRIMT’s real-world applicability on a Kinova Jaco robot.
@@ 2 Related Works and Preliminaries
Foundation models refer to large-scale pre-trained models with strong generalization and reasoning capabilities across tasks [bommasani2021opportunities]. Recent work has explored leveraging FMs to address the reward engineering challenge in RL. One line of research uses coding LLMs to directly generate executable reward functions [maeureka, venutocode, xietext2reward]. Another employs VLMs as contrastive reward signals [ma2023liv, mahmoudieh2022zero, rocamonde2023vision, sontakke2023roboclip]; for example, RoboCLIP [sontakke2023roboclip] rewards agents by aligning trajectory images with task descriptions or demonstrations. However, such explicit FM-based reward signals are often noisy and high-variance [wang2024rl]. Recent approaches have adopted the PbRL paradigm, using FMs as synthetic evaluators to generate trajectory-level preference labels and train reward models: PrefCLM [wang2025prefclm] and RL-SaLLM-F [tu2024online] leverage LLMs to analyze numerical state-action sequences, while RL-VLM-F [wang2024rl] uses VLMs to assess final-state images of robot trajectories. These approaches have shown improved performance over FM-generated scalar rewards. Our work builds on this PbRL-with-FM direction but introduces two key innovations. First, PRIMT adopts a multimodal evaluation scheme that combines VLM and LLM perspectives via hierarchical neuro-symbolic fusion, improving the robustness and quality of synthetic supervision. Second, rather than using FMs solely for passive evaluation, we actively incorporate them to facilitate reward learning via trajectory synthesis.
@@
PbRL aims to learn a reward model $r_{\psi}$ from human comparative feedback over pairs of robot trajectories [lee2021pebble]. A trajectory $\sigma$ is defined as a sequence of states and actions $\{(s_{1},a_{1})\ldots,(s_{T},a_{T})\}$ with a length of $T$. The annotator provides a preference label $\Upsilon\in\{-1,0,1\}$ for each pair $(\sigma^{A},\sigma^{B})$, where $\Upsilon=1$ indicates that $\sigma^{A}$ is preferred, 0 means $\sigma^{B}$ is preferred, and 1-1 denotes indecision. A preference predictor is constructed using the Bradley-Terry model [bradley1952rank] to estimate the preference probabilities. The likelihood that $\sigma^{A}$ is preferred over $\sigma^{B}$ is computed as:
@@
PbRL relies on informative preference queries to train effective reward models. However, trajectory pairs often exhibit low task-relevant diversity, leading to query ambiguity [lee2021b]. This ambiguity is especially pronounced in early training stages, when trajectories are uniformly low-quality and incoherent due to randomly initialized policies [casper2023open]. Prior works address this by selecting maximally distinguishable or uncertain samples [feng2024comparing, biyik2019asking, lee2021b] or initializing the reward model with expert demonstrations [biyik2022learning, palan2019learning]. More recent works employ LLMs to revise ambiguous trajectories during training, assuming that the edited versions are task-complete and preferred [tu2024online]. In contrast, we proactively initialize the replay buffer with LLM-generated trajectories that are diverse and task-aligned. Unlike prior work, we do not assume these are optimal, but use them as preference anchors to support more informative and efficient early-stage evaluation.
@@
Another core challenge in PbRL is the granularity mismatch between trajectory-level preference supervision and the desired state-action-level reward signal [vermahindsight]. This mismatch introduces uncertainty in attributing credit to specific decisions, impairing both the alignment and generalization of the learned reward model [ke2018sparse]. Prior work mitigates this issue by training transformer-based world models that estimate state importance [vermahindsight, kim2022preference, zhao2024prefmmt], or by collecting additional human annotations to highlight key moments [holk2024polite]. In contrast, our approach requires neither extra supervision nor architectural changes. Inspired by causal counterfactual reasoning [mesnard2021counterfactual, pearl2010causal, yang2021top, verma2020counterfactual], we prompt LLMs to generate hindsight-based counterfactual trajectories by minimally editing key decision points in the preferred trajectory to reverse the preference. By asking, “What minimal change would make this trajectory less preferred?”, we obtain contrastive examples that expose the underlying reasons for preference. To effectively leverage these counterfactuals in reward learning, we introduce a causal-aware auxiliary loss. It enforces reward separation at edited points while maintaining consistency in the unedited parts, leading to more precise credit assignment.
@@ 3.1 Multimodal Feedback Generation and Fusion
Given a trajectory pair $(\sigma^{A},\sigma^{B})$ sampled from the trajectory buffer, we first obtain their textual projections $text(\sigma^{A})$ and $text(\sigma^{B})$ for LLM-based evaluation, following [tu2024online]. These projections organize each trajectory into dimension-specific sequences, capturing structured temporal patterns across state and action components in a format that enhances semantic interpretability. For VLM-based evaluation, instead of using all frames or final-state images as in prior work [wang2024rl], we propose a hybrid keyframe extraction method to capture both low-level motion cues and high-level behavior transitions while avoiding visual overload: i) near-zero velocity detection identifies frames where the robot motion is minimal, typically marking subgoal completions or transitional pauses [shridhar2023perceiver]; ii) smoothing residual peaks detect high-curvature or abrupt motion transitions by comparing the raw trajectory to its smoothed version, capturing key motion shifts [akgun2012keyframe]; and iii) change point detection segments the trajectory into semantically coherent phases to identify structural high-level task changes [truong2020selective]. We take the union of the selected frames from all three methods, together with the first and last steps of the trajectory, to form the final keyframe sets $kvis(\sigma^{A})$ and $kvis(\sigma^{B})$. Full details are provided in Appendix B.
@@
To efficiently model these latent dependency structures among inputs, heuristics, and decisions, we employ Probabilistic Soft Logic (PSL) [bach2017hinge], a probabilistic framework representing entities of interest as logical atoms interconnected by weighted first-order logic rules. Specifically, we define four rules to guide inter-modal preference fusion:
@@ 3.2 Bidirectional Trajectory Synthesis
Prior to PbRL training, we employ LLMs to generate bootstrapped trajectories that exhibit diverse, semantically meaningful, and task-aligned behaviors, providing a warm-start initialization for the trajectory buffer. Inspired by structured code-generation paradigms [liang2023code, kwon2024language], we adopt a three-step CoT strategy: i) generate a high-level, multi-step action plan from the task specification; ii) translate each step into executable code snippets that implement concrete motion primitives; and iii) execute these programs under varied initial conditions (e.g., robot start positions) and strategy parameters (e.g., height to approach the target) to synthesize a diverse set of plausible trajectories. Compared to directly prompting LLMs to generate low-level trajectory arrays [tu2024online], our method improves physical feasibility and semantic coherence by grounding trajectory synthesis in program logic. The generated trajectories are considered as bootstrapped demonstrations rather than optimal ones, subsequently evaluated by our multimodal feedback module. Combined with strategic sampling schemes, such as uncertainty-based sampling [lee2021b], these trajectories serve as informative preference anchors when paired with exploration trajectories, reducing ambiguity in early-stage preference queries and accelerating reward learning.
@@
During PbRL training, whenever a clear preference is identified by the multimodal feedback module, we prompt LLMs to perform hindsight reasoning to generate counterfactual variants of the preferred trajectory. This process follows a three-step reasoning pattern based on the structural causal model [pearl2010causal]:
@@
ii) Action: Select a key step $t^{*}\in T^{*}$ for minimal intervention, generating a counterfactual trajectory $\sigma^{*}_{cf}$ that reverses the original preference. This involves modifying some critical state-action features at the selected step, such as introducing a small gripper delay or adding a local end-effector position perturbation. The rest of the trajectory remains identical to the original while we allow the LLMs to apply light smoothing to the immediate neighbors (e.g., 2-3 steps before and after the intervention) to ensure physical continuity and avoid abrupt state transitions. Multiple counterfactual variants can be generated through repeated LLM sampling, providing a diverse set of sub-preferred alternatives. Following the minimal edit principle [goyal2019counterfactual, peng2023diagnosis], we filter the generated counterfactuals based on the L1 distance between the edited state-action pairs and the original, ensuring a small deviation threshold.